OpenStudy (anonymous):

find the perimeter of a triangle with lengths (5x-4) , (3x+5) and (6x-12)

OpenStudy (anonymous):

5x-4+3x+5+6x-12 = 14x-11

OpenStudy (anonymous):

(5x-4)+ (3x+5) + (6x-12) = 14x -11

OpenStudy (anonymous):

5x-4+3x+5+6x-12 5x+3x+6x-4-12 14x-11

OpenStudy (anonymous):

oops there was a mistake -4-12=-16 so,14x-16=180+16 14x=196/14 x=14